Angel06:

In a race , Ram covers 5 km in 20 min. How much distance will he cover in 100 min ? A.25 km B.26 km C.35 km D.40 km

JamesTDG:

This is some simple math. 100/20=10/2=5 | 5*5=25 | Your answer is A. 25